OpenStudy (anonymous):

I need help with the following things: Write the equation of the line that passes through (2, 2) and (6, 3) in standard form. Write the equation of the line that passes through (3, 6) and (0, –3) in slope-intercept form Write the equation of the line that passes through (2, 5) with a slope of 3 in point-slope form. Write the equation of the line that passes through (–1, 5) with a slope of 2 in standard form

OpenStudy (anonymous):

\[+7\]

OpenStudy (anonymous):

Huh?

OpenStudy (anonymous):

general equation y2-y1= m(x2-x1) since you have 2 points, determine the slope, which is m once you have m, substitute a point and m into the same equation y-y1=m(x-x1) and then do whatever is required in order to get the form Ax+By=C for point intercept form y=mx+b once you have m substitute a point and m into the equation, then solve for b then rewrite it only substituting m and b

OpenStudy (anonymous):

Can you show me one?

OpenStudy (anonymous):

alright then, lets start with the first problem we have the points (2 ,2) & (6,3) so we substitute those into the the point slope form given y2-y1=m(x2-x1) and find m 3-2 = m(6-2) 1 = m *4 m= 1/4 once we have m we substitute this back into the equation with one of the two points so y-2= (1/4) (x-2), at this point, you multiply, distribute and do whatever you need to do to this system in order to make it look like Ax+Bx=C or if you need it in point intercept form once you have m, substitute a point and m into the point intercept equation and solve for b y=mx+b remember that x intercept is always (x,0) and y- intecept is always (0,y)

OpenStudy (anonymous):

Is this the answers to my questions?? Lol

OpenStudy (anonymous):

looks like only for the first one

OpenStudy (anonymous):

ya its only for the first one

OpenStudy (anonymous):

yes it is the answer lyssa

OpenStudy (anonymous):

I know but only for the first one lol